# SoK: Cybersecurity Assessment of Humanoid Ecosystem

### v1

@@ 4.1 Holistic Humanoid Threat Landscape
Attack ID Attack Vector Description Ref. Physical Layer - Hardware Integrity and Energy Manipulation P-A1 Firmware reflashing Quarta et al. demonstrated that the FlexPendant boot image delivered by the controller lacks code signing, enabling installation of a malicious firmware that persists across reboot; they used this to falsify the User Interface (UI), misleading the operator about robot state. Humanoids mostly use unsigned firmware images and weak update channels. [76] P-A2 Sensor bias via local fields Many studies [89, 14, 88] have empirically demonstrated sensor-bias attacks at the modality level (e.g., magnetic/Hall encoders, LiDAR, MEMS IMUs, microphones) and are platform-agnostic for robots that deploy the same sensor type. [89] P-A3 Power rail manipulation Zhan et al. demonstrated that attackers can manipulate Qi-compliant wireless chargers by modulating their input supply, and control the charging systems to cause overheating and overcharging [121]. humanoids with external DC adapters, can force unsafe charging conditions, potentially triggering system resets that disrupt critical control loops. [121] P-A4 Field tampering Belikovetsky et al. weakened a 3D-printed propeller via CAD/print-path manipulation, which later fractured mid-flight and caused a crash; the defect escaped visual inspection [7]. This research describes how supply chain attacks on suppliers of hardware components of robotics can serve as silent vectors for physical compromise. [7] P-A5 Acoustic side-channels Shah et al. showed that smartphone microphones can fingerprint robot motions and reconstruct workflows from gearbox acoustics, leaking task-level behavior [84]. LidarPhone used robot vacuum’s LiDAR as a laser mic to eavesdrop on speech via vibrations of nearby objects. humanoids with similar gearboxes/LiDAR, are vulnerable to similar exploitation [81]. [81, 84] Sensing and Perception Layer - Sensor Fidelity, Privacy, and Signal Integrity SP-A1 Active optical spoofing Shin et al. demonstrated that injecting timed infrared pulses into VLP-16 LiDAR systems can create phantom detection points or blind sectors [88]. Cao et al. achieved a 92.7% success rate in removing approximately 90% of target obstacles from point clouds, effectively hiding moving pedestrians from detection. These perception-layer attacks pose threat for humanoid robots using similar LiDAR systems. [88, 14] SP-A2 Sensor blinding and jamming Petit et al. experimentally blinded cameras with high-intensity light and spoofed LiDAR with laser pulses, producing false obstacles or loss of returns. On humanoids, it can mislead footstep planning at edges and cause falls [69]. [69] SP-A3 Resonant ultrasound and EMI injection Son et al. destabilized drones by driving MEMS gyroscopes at acoustic resonance [91]; Trippel et al. injected signals into MEMS accelerometers via sound to control their digital outputs [99]. On humanoids, similar IMU bias can corrupt state estimation and cascade to balance control, risking falls. [91, 99] SP-A4 Physical adversarial patches Thys et al. printed patches that hide persons from YOLOv2, markedly reducing detections in real-world video [98]; Eykholt et al. crafted sticker attacks that fooled Faster R-CNN in 85.9% of frames in lab tests (40.2% outdoor) [24]. On humanoids, such perception-layer evasion can conceal humans/obstacles and propagate errors to planning and navigation. [98, 24] SP-A5 Sensor-based privacy leakage Giaretta et al. found that Pepper’s HTTP interfaces and unauthenticated endpoints expose camera access with plaintext credentials [27], Campazas-Vega et al. observed similar unencrupted HTTP-only communication causing data leak [102]. [27, 102] Data Processing Layer - Runtime Pipelines, Memory safety, and Control Loop Consistency DP-A1 Timing channel interference Li et al. showed that microsecond-scale contention on PREEMPT-RT (real-time Linux) causes deadline misses, demonstrated when a humanoid research platform ROBOTIS OP3’s 1 kHz balance loop failed and caused the robot to fall [43]. Proctor measured similar microsecond-level MCU jitter from cache and bus contention, confirming that timing interference can destabilize torque control loops [75]. [43, 75] DP-A2 Memory safety exploits CVE-2018-5268 documented how crafted JPEG2000 images trigger heap-based buffer overflows in OpenCV 3.3.1, compromising perception pipelines that process untrusted visual inputs [17]. Alias Robotics identified 86 vulnerabilities in Universal Robots controllers, including buffer overflows in real-time joint control code, demonstrating how data-processing layer attacks can cascade to planning and actuation systems in humanoid robots [79]. [17, 79] DP-A3 Race conditions and TOCTOU flaws ROS 2 documented a MultiThreadedExecutor race condition where messages are dequeued before publishers complete their writes, exposing stale or partial data (rclcpp #1007). This can propagate stale odometry from data-processing to planning and control systems in humanoid robots [97]. [97] DP-A4 Estimator bias tampering Shen et al. showed GPS spoofing can seize multi sensor fusion (MSF) Kalman Filter (KF) localization and induce >10 m of lateral error before fault alarms fired, letting estimator bias cascade into planning and control [85], while Cao et al. removed obstacles from LiDAR and LiDAR–camera fusion so detectors missed objects in closed-track tests [14]. [14] Middleware Layer - Secure Communications, Naming, and Orchestration MW-A1 Topic spoofing and participant impersonation Mayoral et al. showed that spoofed ROS publishers/subscribers let attackers inject commands and seize endpoints, escalating a middleware breach into planning and actuation effects [106]. Di Luoffo et al. showed credential-masquerading attacks against DDS/SROS2 that allow participant impersonation and message forgery, with unsecured deployments further exposing RTPS discovery identifiers that facilitate spoofing [21]. [106, 21] MW-A2 Replay and stale frame injection Lauser et al. identified a replay vulnerability in DDS-Security’s encryption protocol that enables stale message injection even with security enabled [40]. In humanoids, replayed odometry and IMU data would propagate through localization and balance systems, compromising stability. [40] MW-A3 Plaintext sniffing and data exfiltration Giaretta et al. demonstrated that SoftBank’s Pepper robot transmitted unencrypted RGB camera and audio streams over Wi-Fi, enabling interception of live sensor data and exposure of personally identifiable information (PII) [27]. [27] MW-A4 Discovery abuse and denial-of-service (DoS) Cibrario et al. demonstrated SlowROS, a low rate DoS attack against ROS Master that exploits missing connection-closure timeouts and causes system-wide ROS disruptions. Humanoids rely on ROS middleware for control topics such as /joint_angles and /imu, a comparable flood could stall message delivery and destabilize balance controllers [9]. [9] Decision-Making Layer - Compromise of Autonomy and Control Logic DM-A1 Adversarial Examples in Perception and Policy Chen et al. showed that a 5 cm adversarial patch inflated visual SLAM error by 7.3× and caused RGB-D tracking loss within 20 frames [15]. Buddareddygari et al. demonstrated that adversarial roadside textures misled deep RL driving policies, inducing emergency stops or wrong-way routing within seconds [13]. In humanoids, such patches could misclassify gestures or inject phantom obstacles, undermining decision making. [15, 13] DM-A2 Sequential perturbation of control policies Shi et al. demonstrated that injecting low-magnitude, time-correlated perturbations into IMU readings destabilized the ANYmal C quadruped, causing falls in 87% of trials despite model predictive control fallbacks. The attack gradually corrupted state estimates across sequential control cycles [86]. Humanoids fusing joint proprioception with inertial data face similar gradual destabilization risks. [86] DM-A3 Concept drift and sensor fusion spoofing Shen et al. introduced FusionRipper, a GPS spoofing attack that stealthily biases multi-sensor fusion localization without triggering health alarms. On Baidu’s Apollo, it laterally shifted the MSF output by over 10 m within minutes while reporting OK status. Humanoids using visual-inertial odometry, ultra wide band (UWB) beacons, or floor-marker fusion may similarly be misled via staged bias injections that exploit KF latency. [85]
@@
Attack ID Attack Vector Description Ref. DM-A4 Reward hacking and specification gaming Krakovna et al. catalogued specification-gaming cases in reinforcement learning, such as CoastRunners agents repeatedly hitting bonus targets instead of finishing races. Domestic humanoid rewarded for staying active near its charging dock might simply spin in place to maximize reward, rather than performing its intended household duties [37]. [37] DM-A5 Model poisoning and physical backdoors Wang et al. proposed TrojanRobot, a backdoor attack that embeds visual triggers into vision-language model (VLM) manipulation policies. Testing on a UR3e robotic arm showed that mundane objects stealthily alter action outputs while maintaining normal performance otherwise [110]. Humanoids using similar vision-language-action pipelines for grasping or tool use face comparable risks of stealthy policy hijacking. [110] DM-A6 Finite State Machine (FSM) corruption and logic bombs Govil et al. introduced ladder logic bombs (LLBs)—malicious code embedded in PLC programs that remains dormant until triggered, then overwrites control states or sensor outputs while evading detection [28]. In humanoids, similar latent attacks could hide within gait or balance controllers, activating during specific pose transitions or sensor events. [28] DM-A7 Control parameter tampering Trend Micro’s Rogue Robots study demonstrated that subtle tampering with ABB IRB 140 control parameters—such as servo gains or sensor offsets—introduces undetected motion deviations that compromise safety [77]. In humanoids, manipulation of joint limits or kinematic parameters could degrade motion safety without triggering controller alarms. [77] DM-A8 Planning logic manipulation Vemprala et al. demonstrated that adversaries can manipulate optimization-based motion planners to produce unsafe trajectories by perturbing cost functions [103]. Since humanoids rely on cost-map planning, similar tampering through false sensor inputs or corrupted avoidance rules could misdirect navigation and compromise safety. [103] DM-A9 Malicious OTA updates NHTSA classified firmware OTA updates as high-impact threat vector, since compromised OTA servers can silently corrupt entire system before detection. Humanoids face similar risks of stealthy sabotage that bypasses safety checks [101]. [101] DM-A10 Federated learning poisoning Bagdasaryan et al. demonstrated that a single malicious participant in federated learning can replace the global model with a backdoored version, preserving accuracy while achieving 100% success on hidden malicious objectives [5]. This vulnerability could enable stealthy manipulation of decision-making policies across multiple deployed humanoid units. [5] DM-A11 Trojaned pretrained models and supply chain threats Wang et al. embedded a backdoor into a pretrained VLM controlling a UR3e robot arm. The model operated normally until visual triggers activated attacker-specified manipulation routines, demonstrating that trojaned models can survive deployment undetected [110]. Humanoid systems using third-party or cloud-hosted models face similar supply-chain risks of malicious behaviors that activate only under specific conditions. [110] Application Layer - Task Scripts, APIs, and OTA Operations AP-A1 Script and state-machine corruption Quarta et al. demonstrated that modifying ABB RAPID control code silently alters robot behavior while status displays showed normal operation [76]. Pogliani et al. showed that URControl robot’s network interface accepted malicious URScript commands, enabling remote code execution with root privileges. In humanoids, such attacks could modify the behavior trees to alter movement patterns or disable safety checks [72]. [76, 72] AP-A2 Unauthenticated APIs and remote override DeMarinis et al. found over 100 ROS systems exposed on the public internet, enabling unauthenticated remote control of robots through basic commands like /cmd_vel [18]. Giaretta et al. discovered that Pepper’s NAOqi API lacked authentication, allowing attackers to access cameras, microphones, and disable safety functions [27]. [18, 27] AP-A3 CI/CD supply chain compromise Oyeronke analyzed firmware supply chains in CI/CD pipelines and showed that attackers could tamper with dependencies, inject code during builds, or steal signing keys to distribute malicious images as legitimate OTA updates [62]. In humanoids, such a poisoned pipeline could enable the silent deployment of trojaned firmware across fleets of robots. [62] AP-A4 Sim-to-real discrepancy and digital-twin overload Xu et al. showed that a single malicious node can saturate cloud multi-robot services, causing 35% packet loss and reducing robot velocity to 0.04-0.05 m/s, leading to mission failures and crashes [113]. This overload disrupts digital twin synchronization: delayed updates create twin-physical misalignment, causing plans to be computed on stale state data [94]. Since latency drives the sim-to-real gap, these synchronization delays result in degraded or unsafe behaviors when control policies expecting timely feedback execute on outdated measurements [95]. [113, 95, 94] AP-A5 Runtime parameter tampering Quarta et al. demonstrated that modifying controller parameters degraded accuracy and induced instability in ABB IRB140 robot, with configuration changes propagating from application to servo levels [76]. For humanoids, similar attacks on ROS-hosted gait parameters (YAML/URDF) could silently destabilize walking while bypassing high-level safety checks. [76] Social Interface Layer - Speech, Gesture and the Human-Trust Channel SI-A1 Inaudible and line-of-sight command injection Zhang et al. showed that ultrasound-modulated, inaudible commands can activate and control voice assistants at distances [122]. Sugawara et al. showed that amplitude-modulated lasers inject commands through closed glass [93]. This enables silent takeover of social interfaces in humanoids that can cascade to decision and application layers. [122, 93] SI-A2 Synthetic or hidden audio Yuan et al. embedded automatic speech recognition (ASR) commands into songs [119]. Li et al. crafted adversarial music that hides ghost commands which robot dialogue systems’ ASR models execute with high success rates while sounding normal to humans [46]. Humanoids working on voice commands can be exploited through social interfaces with effects cascading to decision/application layers. [119, 46] SI-A3 Visual spoofing and adversarial cues Thys et al. demonstrated that printed adversarial patches reduced YOLOv2 person detection allowing 61-74% of people to bypass detection [98]. Wu et al. used projected structured-light patterns from 2D photos to spoof commercial 3D face authentication with 79.4% success (depth-only) [111]. For humanoids, adversarial clothing can enable user impersonation. [98, 111] SI-A4 Covert A/V eaves-dropping via social channels Giaretta et al. found that Pepper’s web interface exposed credentials, enabling privilege escalation for covert surveillance [27]. IOActive demonstrated remote access to cameras and microphones on Pepper/NAO and UBTECH Alpha-2 robots [61]. Denning et al. showed that consumer telepresence robots could be compromised to stream live video for home surveillance [20]. Such compromises in humanoids enables covert surveillance and intrusion of privacy. [27, 20, 61] SI-A5 Trust manipulation and social engineering Aroyo et al. used iCub robot in social-engineering scenarios where participants disclosed sensitive data and were induced to gamble their winnings, demonstrating overtrust-driven risk-taking. Wolfert et al. showed Pepper could tailgate into secure areas and convinced all participants to insert unknown USB devices. Such compromise can enable phishing attacks and coercion of unsafe user actions. [8, 4]

### v2

@@ 4.1 Holistic Humanoid Threat Landscape
Attack ID Attack Vector Description Ref. Physical Layer - Hardware Integrity and Energy Manipulation P-A1 Firmware reflashing Quarta et al. demonstrated that the FlexPendant boot image delivered by the controller lacks code signing, enabling installation of a malicious firmware that persists across reboot; they used this to falsify the User Interface (UI), misleading the operator about robot state. Humanoids mostly use unsigned firmware images and weak update channels. [81] P-A2 Sensor bias via local fields Many studies [90, 14, 89] have empirically demonstrated sensor-bias attacks at the modality level (e.g., magnetic/Hall encoders, LiDAR, MEMS IMUs, microphones) and are platform-agnostic for robots that deploy the same sensor type. [90] P-A3 Power rail manipulation Zhan et al. demonstrated that attackers can manipulate Qi-compliant wireless chargers by modulating their input supply, and control the charging systems to cause overheating and overcharging [121]. humanoids with external DC adapters, can force unsafe charging conditions, potentially triggering system resets that disrupt critical control loops. [121] P-A4 Field tampering Belikovetsky et al. weakened a 3D-printed propeller via CAD/print-path manipulation, which later fractured mid-flight and caused a crash; the defect escaped visual inspection [7]. This research describes how supply chain attacks on suppliers of hardware components of robotics can serve as silent vectors for physical compromise. [7] P-A5 Acoustic side-channels Shah et al. showed that smartphone microphones can fingerprint robot motions and reconstruct workflows from gearbox acoustics, leaking task-level behavior [85]. LidarPhone used robot vacuum’s LiDAR as a laser mic to eavesdrop on speech via vibrations of nearby objects. humanoids with similar gearboxes/LiDAR, are vulnerable to similar exploitation [83]. [83, 85] Sensing and Perception Layer - Sensor Fidelity, Privacy, and Signal Integrity SP-A1 Active optical spoofing Shin et al. demonstrated that injecting timed infrared pulses into VLP-16 LiDAR systems can create phantom detection points or blind sectors [89]. Cao et al. achieved a 92.7% success rate in removing approximately 90% of target obstacles from point clouds, effectively hiding moving pedestrians from detection. These perception-layer attacks pose threat for humanoid robots using similar LiDAR systems. [89, 14] SP-A2 Sensor blinding and jamming Petit et al. experimentally blinded cameras with high-intensity light and spoofed LiDAR with laser pulses, producing false obstacles or loss of returns. On humanoids, it can mislead footstep planning at edges and cause falls [74]. [74] SP-A3 Resonant ultrasound and EMI injection Son et al. destabilized drones by driving MEMS gyroscopes at acoustic resonance [92]; Trippel et al. injected signals into MEMS accelerometers via sound to control their digital outputs [103]. On humanoids, similar IMU bias can corrupt state estimation and cascade to balance control, risking falls. [92, 103] SP-A4 Physical adversarial patches Thys et al. printed patches that hide persons from YOLOv2, markedly reducing detections in real-world video [102]; Eykholt et al. crafted sticker attacks that fooled Faster R-CNN in 85.9% of frames in lab tests (40.2% outdoor) [93]. On humanoids, such perception-layer evasion can conceal humans/obstacles and propagate errors to planning and navigation. [102, 93] SP-A5 Sensor-based privacy leakage Giaretta et al. found that Pepper’s HTTP interfaces and unauthenticated endpoints expose camera access with plaintext credentials [25], Campazas-Vega et al. observed similar unencrypted HTTP-only communication causing data leak [12]. [25, 12] Data Processing Layer - Runtime Pipelines, Memory safety, and Control Loop Consistency DP-A1 Timing channel interference Li et al. showed that microsecond-scale contention on PREEMPT-RT (real-time Linux) causes deadline misses, demonstrated when a humanoid research platform ROBOTIS OP3’s 1 kHz balance loop failed and caused the robot to fall [45]. Proctor measured similar microsecond-level MCU jitter from cache and bus contention, confirming that timing interference can destabilize torque control loops [80]. [45, 80] DP-A2 Memory safety exploits CVE-2018-5268 documented how crafted JPEG2000 images trigger heap-based buffer overflows in OpenCV 3.3.1, compromising perception pipelines that process untrusted visual inputs [61]. Alias Robotics identified 86 vulnerabilities in Universal Robots controllers, including buffer overflows in real-time joint control code, demonstrating how data-processing layer attacks can cascade to planning and actuation systems in humanoid robots [26]. [61, 26] DP-A3 Race conditions and TOCTOU flaws ROS 2 documented a MultiThreadedExecutor race condition where messages are dequeued before publishers complete their writes, exposing stale or partial data (rclcpp #1007). This can propagate stale odometry from data-processing to planning and control systems in humanoid robots [101]. [101] DP-A4 Estimator bias tampering Shen et al. showed GPS spoofing can seize multi sensor fusion (MSF) Kalman Filter (KF) localization and induce >10 m of lateral error before fault alarms fired, letting estimator bias cascade into planning and control [86], while Cao et al. removed obstacles from LiDAR and LiDAR–camera fusion so detectors missed objects in closed-track tests [14]. [14] Middleware Layer - Secure Communications, Naming, and Orchestration MW-A1 Topic spoofing and participant impersonation Mayoral et al. showed that spoofed ROS publishers/subscribers let attackers inject commands and seize endpoints, escalating a middleware breach into planning and actuation effects [57]. Di Luoffo et al. showed credential-masquerading attacks against DDS/SROS2 that allow participant impersonation and message forgery, with unsecured deployments further exposing RTPS discovery identifiers that facilitate spoofing [21]. [57, 21] MW-A2 Replay and stale frame injection Lauser et al. identified a replay vulnerability in DDS-Security’s encryption protocol that enables stale message injection even with security enabled [44]. In humanoids, replayed odometry and IMU data would propagate through localization and balance systems, compromising stability. [44] MW-A3 Plaintext sniffing and data exfiltration Giaretta et al. demonstrated that SoftBank’s Pepper robot transmitted unencrypted RGB camera and audio streams over Wi-Fi, enabling interception of live sensor data and exposure of personally identifiable information (PII) [25]. [25] MW-A4 Discovery abuse and denial-of-service (DoS) Cibrario et al. demonstrated SlowROS, a low rate DoS attack against ROS Master that exploits missing connection-closure timeouts and causes system-wide ROS disruptions. Humanoids rely on ROS middleware for control topics such as /joint_angles and /imu, a comparable flood could stall message delivery and destabilize balance controllers [16]. [16] Decision-Making Layer - Compromise of Autonomy and Control Logic DM-A1 Adversarial Examples in Perception and Policy Chen et al. showed that a 5 cm adversarial patch inflated visual SLAM error by 7.3× and caused RGB-D tracking loss within 20 frames [15]. Buddareddygari et al. demonstrated that adversarial roadside textures misled deep RL driving policies, inducing emergency stops or wrong-way routing within seconds [11]. In humanoids, such patches could misclassify gestures or inject phantom obstacles, undermining decision making. [15, 11] DM-A2 Sequential perturbation of control policies Shi et al. demonstrated that injecting low-magnitude, time-correlated perturbations into IMU readings destabilized the ANYmal C quadruped, causing falls in 87% of trials despite model predictive control fallbacks. The attack gradually corrupted state estimates across sequential control cycles [87]. Humanoids fusing joint proprioception with inertial data face similar gradual destabilization risks. [87]
@@
Attack ID Attack Vector Description Ref. DM-A3 Concept drift and sensor fusion spoofing Shen et al. introduced FusionRipper, a GPS spoofing attack that stealthily biases multi-sensor fusion localization without triggering health alarms. On Baidu’s Apollo, it laterally shifted the MSF output by over 10 m within minutes while reporting OK status. Humanoids using visual-inertial odometry, ultra wide band (UWB) beacons, or floor-marker fusion may similarly be misled via staged bias injections that exploit KF latency. [86] DM-A4 Reward hacking and specification gaming Krakovna et al. catalogued specification-gaming cases in reinforcement learning, such as CoastRunners agents repeatedly hitting bonus targets instead of finishing races. Domestic humanoid rewarded for staying active near its charging dock might simply spin in place to maximize reward, rather than performing its intended household duties [41]. [41] DM-A5 Model poisoning and physical backdoors Wang et al. proposed TrojanRobot, a backdoor attack that embeds visual triggers into vision-language model (VLM) manipulation policies. Testing on a UR3e robotic arm showed that mundane objects stealthily alter action outputs while maintaining normal performance otherwise [110]. Humanoids using similar vision-language-action pipelines for grasping or tool use face comparable risks of stealthy policy hijacking. [110] DM-A6 Finite State Machine (FSM) corruption and logic bombs Govil et al. introduced ladder logic bombs (LLBs)-malicious code embedded in PLC programs that remains dormant until triggered, then overwrites control states or sensor outputs while evading detection [27]. In humanoids, similar latent attacks could hide within gait or balance controllers, activating during specific pose transitions or sensor events. [27] DM-A7 Control parameter tampering Trend Micro’s Rogue Robots study demonstrated that subtle tampering with ABB IRB 140 control parameters-such as servo gains or sensor offsets-introduces undetected motion deviations that compromise safety [52]. In humanoids, manipulation of joint limits or kinematic parameters could degrade motion safety without triggering controller alarms. [52] DM-A8 Planning logic manipulation Vemprala et al. demonstrated that adversaries can manipulate optimization-based motion planners to produce unsafe trajectories by perturbing cost functions [104]. Since humanoids rely on cost-map planning, similar tampering through false sensor inputs or corrupted avoidance rules could misdirect navigation and compromise safety. [104] DM-A9 Malicious OTA updates NHTSA classified firmware OTA updates as high-impact threat vector, since compromised OTA servers can silently corrupt entire system before detection. Humanoids face similar risks of stealthy sabotage that bypasses safety checks [63]. [63] DM-A10 Federated learning poisoning Bagdasaryan et al. demonstrated that a single malicious participant in federated learning can replace the global model with a backdoored version, preserving accuracy while achieving 100% success on hidden malicious objectives [5]. This vulnerability could enable stealthy manipulation of decision-making policies across multiple deployed humanoid units. [5] DM-A11 Trojaned pretrained models and supply chain threats Wang et al. embedded a backdoor into a pretrained VLM controlling a UR3e robot arm. The model operated normally until visual triggers activated attacker-specified manipulation routines, demonstrating that trojaned models can survive deployment undetected [110]. Humanoid systems using third-party or cloud-hosted models face similar supply-chain risks of malicious behaviors that activate only under specific conditions. [110] Application Layer - Task Scripts, APIs, and OTA Operations AP-A1 Script and state-machine corruption Quarta et al. demonstrated that modifying ABB RAPID control code silently alters robot behavior while status displays showed normal operation [81]. Pogliani et al. showed that URControl robot’s network interface accepted malicious URScript commands, enabling remote code execution with root privileges. In humanoids, such attacks could modify the behavior trees to alter movement patterns or disable safety checks [77]. [81, 77] AP-A2 Unauthenticated APIs and remote override DeMarinis et al. found over 100 ROS systems exposed on the public internet, enabling unauthenticated remote control of robots through basic commands like /cmd_vel [18]. Giaretta et al. discovered that Pepper’s NAOqi API lacked authentication, allowing attackers to access cameras, microphones, and disable safety functions [25]. [18, 25] AP-A3 CI/CD supply chain compromise Oyeronke analyzed firmware supply chains in CI/CD pipelines and showed that attackers could tamper with dependencies, inject code during builds, or steal signing keys to distribute malicious images as legitimate OTA updates [56]. In humanoids, such a poisoned pipeline could enable the silent deployment of trojaned firmware across fleets of robots. [56] AP-A4 Sim-to-real discrepancy and digital-twin overload Xu et al. showed that a single malicious node can saturate cloud multi-robot services, causing 35% packet loss and reducing robot velocity to 0.04-0.05 m/s, leading to mission failures and crashes [114]. This overload disrupts digital twin synchronization: delayed updates create twin-physical misalignment, causing plans to be computed on stale state data [98]. Since latency drives the sim-to-real gap, these synchronization delays result in degraded or unsafe behaviors when control policies expecting timely feedback execute on outdated measurements [99]. [114, 99, 98] AP-A5 Runtime parameter tampering Quarta et al. demonstrated that modifying controller parameters degraded accuracy and induced instability in ABB IRB140 robot, with configuration changes propagating from application to servo levels [81]. For humanoids, similar attacks on ROS-hosted gait parameters (YAML/URDF) could silently destabilize walking while bypassing high-level safety checks. [81] Social Interface Layer - Speech, Gesture and the Human-Trust Channel SI-A1 Inaudible and line-of-sight command injection Zhang et al. showed that ultrasound-modulated, inaudible commands can activate and control voice assistants at distances [123]. Sugawara et al. showed that amplitude-modulated lasers inject commands through closed glass [96]. This enables silent takeover of social interfaces in humanoids that can cascade to decision and application layers. [123, 96] SI-A2 Synthetic or hidden audio Yuan et al. embedded automatic speech recognition (ASR) commands into songs [120]. Li et al. crafted adversarial music that hides ghost commands which robot dialogue systems’ ASR models execute with high success rates while sounding normal to humans [46]. Humanoids working on voice commands can be exploited through social interfaces with effects cascading to decision/application layers. [120, 46] SI-A3 Visual spoofing and adversarial cues Thys et al. demonstrated that printed adversarial patches reduced YOLOv2 person detection allowing 61-74% of people to bypass detection [102]. Wu et al. used projected structured-light patterns from 2D photos to spoof commercial 3D face authentication with 79.4% success (depth-only) [112]. For humanoids, adversarial clothing can enable user impersonation. [102, 112] SI-A4 Covert A/V eaves-dropping via social channels Giaretta et al. found that Pepper’s web interface exposed credentials, enabling privilege escalation for covert surveillance [25]. IOActive demonstrated remote access to cameras and microphones on Pepper/NAO and UBTECH Alpha-2 robots [67]. Denning et al. showed that consumer telepresence robots could be compromised to stream live video for home surveillance [20]. Such compromises in humanoids enables covert surveillance and intrusion of privacy. [25, 20, 67] SI-A5 Trust manipulation and social engineering Aroyo et al. used iCub robot in social-engineering scenarios where participants disclosed sensitive data and were induced to gamble their winnings, demonstrating overtrust-driven risk-taking. Wolfert et al. showed Pepper could tailgate into secure areas and convinced all participants to insert unknown USB devices. Such compromise can enable phishing attacks and coercion of unsafe user actions. [111, 4]